OpenStudy (anonymous):

write a pseudocode or draw a flow chart that calculates the discount on a given sale.......assuming that the discount rate is dependent on the quantity of the items sold and is determined.....0-10=o percent....11-24=10 percent....25-50=15 percent......51 or more=20 percent....

OpenStudy (anonymous):

hope this makes sense

OpenStudy (anonymous):

int quantity float sale, discount, unit_price print "enter the sale?" //calculating the quantity quantity = sale / unit_price //calculating the discounts if ( quantity<=10 ) print "Discount 0" else if (quantity > 10 and quantity<25) discount = sale *0.1 else if(quantity >= 25 and quantity<=50) discount=sale*0.15 else discount=sale*0.2

OpenStudy (anonymous):

thanks alot 4 trying 2 answer ma question and making ma work easier....god bless u....